OpenStudy (anonymous):

Solve for x: ln(2x+1)-lnx=2ln3

OpenStudy (anonymous):

start with \[\ln\left(\frac{2x+1}{x}\right)=\ln(9)\] and then solve \[\frac{2x+1}{x}=9\] for \(x\)

OpenStudy (anonymous):

Thanks so much, I got it!

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."